### v1

@@ Abstract
Time series anomaly detection (TSAD) finds many applications such as monitoring environmental sensors, industry KPIs, patient biomarkers, etc. A two-fold challenge for TSAD is a versatile and unsupervised model that can detect various different types of time series anomalies (spikes, discontinuities, trend shifts, etc.) without any labeled data. Modern neural networks have outstanding ability in modeling complex time series. Self-supervised models in particular tackle unsupervised TSAD by transforming the input via various augmentations to create pseudo anomalies for training. However, their performance is sensitive to the choice of augmentation, which is hard to choose in practice, while there exists no effort in the literature on data augmentation tuning for TSAD without labels. Our work aims to fill this gap. We introduce TSAP for TSA “on autoPilot”, which can (self-)tune augmentation hyperparameters end-to-end. It stands on two key components: a differentiable augmentation architecture and an unsupervised validation loss to effectively assess the alignment between augmentation type and anomaly type. Case studies show TSAP’s ability to effectively select the (discrete) augmentation type and associated (continuous) hyperparameters. In turn, it outperforms established baselines, including SOTA self-supervised models, on diverse TSAD tasks exhibiting different anomaly types.
@@ 1 Introduction
Anomaly detection (AD) is a critical task in various domains such as cybersecurity, healthcare, and finance. AD is especially important in time series to ensure system safety and reliability. Thus, there exists a large body of work on time series AD as presented in various surveys [5, 15]. Recent progress on self-supervised learning (SSL) has changed the field of AD, offering significant improvements over traditional unsupervised (or one-class) learning approaches. The main advantage of SSL lies in its ability to self-generate labeled samples, i.e., pseudo anomalies, within the input space. This enables a more focused exploration of a plausible subspace based on the semantics reflected in the pseudo anomalies, rather than an exhaustive, impractical search of the entire space.
At the heart of SSL-based AD are data augmentation functions. These functions are used to create pseudo labels for the (self-)supervised training of an anomaly detector, such as by predicting whether the input is augmented or not [20], which augmentation function is used [14], or using contrastive learning [31]. In all these approaches, the success of SSL-based AD highly depends on the degree to which the augmented data mimics the true anomalies, as outlined in [35].
There exist few approaches for tuning data augmentation functions without labels, however, they have limitations. Some rely on non-differentiable validation losses [36], which are unsuitable for end-to-end learning frameworks. Others only address continuous hyperparameters, neglecting discrete ones, which are left for manual adjustment [34]. For instance, in the image domain, the CutOut augmentation cannot mimic semantic class anomalies (e.g. cats vs. cars), no matter how one tunes its (continuous) hyperparameters [34], mainly due to the mismatch in the discrete choice (i.e., the augmentation type). Moreover, none of these efforts addresses anomaly detection for time series data, which is the focus of our work.
In this paper we introduce TSAP, a novel approach for SSL-based time series anomaly detection (TSAD) “on autoPilot” equipped with end-to-end hyperparameter tuning. Effectively tuning both discrete and continuous hyperparameters of augmentation enables TSAP to be an automated anomaly detector that is most suitable for a given task. TSAP stands on two main components: ($i$) a differentiable parameterized augmentation model and ($ii$) an unsupervised validation loss that measures the alignment between the augmented data and the unlabeled test data, quantifying the extent to which the former mimics the true anomalies. We summarize our main contributions as follows:
@@
    Problem: Our work is the first attempt to tune both discrete and continuous hyperparameters of data augmentation in SSL-based TSAD, without any labels at training time.
@@
    New TSAD Method: We propose TSAP  [^1] The code will be published once the paper is accepted., which accommodates a list of various time series anomaly types and enables automatic tuning of related hyperparameters (magnitude, duration, etc.) by a differentiable validation loss that quantifies alignment of augmented data and unlabeled test data.
@@
    Effectiveness: By carefully selecting augmentation type and its (continuous) hyperparameters, our self-tuning TSAP outperforms existing unsupervised and self-supervised approaches, including the SOTA NeuTraL-AD [26] which also employs learnable augmentations, across diverse TSAD tasks.
2 Preliminaries
2.1 Time Series Anomaly Detection
This work focuses on time series anomaly detection. Consider a univariate time series $\mathbf{x}=\{x_{1},x_{2},\ldots,x_{K}\}$, where $\mathbf{x}$ is a sequentially ordered collection of $K$ data points. Each $x_{k}\in\mathbb{R}$ corresponds to a scalar observation at each time step. Then, let $\mathcal{D}_{\mathrm{trn}}$ be a set of training data consisting of only normal time series, and $\mathcal{D}_{\mathrm{test}}$ be a set of unlabeled test data containing both normal and anomalous time series. Given $\mathcal{D}_{\mathrm{trn}}$ and $\mathcal{D}_{\mathrm{test}}$, the problem is to detect all anomalous time series in $\mathcal{D}_{\mathrm{test}}$, i.e., to assign the accurate label $y\in\{-1,+1\}$ for each $\mathbf{x}\in\mathcal{D}_{\mathrm{test}}$.
2.2 Self-supervised Anomaly Detectors
Let $f_{\mathrm{aug}}(\cdot;\boldsymbol{\phi})$ be a data augmentation function, such as rotation or cut-out in the image domain (a discrete choice), that outputs pseudo anomalies. A given $f_{\mathrm{aug}}$ also exhibits continuous hyperparameter(s), such as angle (for rotation) or width and height (for cutout). Then, a popular approach [20] is to train a detector model $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta})$ to classify between normal data $\mathcal{D}_{\mathrm{trn}}$ and augmented data $\mathcal{D}_{\mathrm{aug}}=\{f_{\mathrm{aug}}(\mathbf{x})\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$. The working assumption is for $f_{\mathrm{det}}$ to predict the (unknown) anomalies in test data as augmented. As such, performance relies on how well the choice of augmentation type (discrete) as well as the choice(s) of its (continuous) hyperparameter(s) mimic the anomalies in $\mathcal{D}_{\mathrm{test}}$. There are various other ways of using SSL for anomaly detection, based on the specific objective function and how the pseudo labels are generated [14, 31], which are similar to each other in principle.
2.3 Data Augmentation on Time Series
To apply SSL to time series data, we need an augmentation function specifically designed for time series. For a given time series $\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}$, $f_{\mathrm{aug}}(\mathbf{x};\mathbf{a})$ transforms $\mathbf{x}$ based on hyperparameters $\mathbf{a}\in\mathcal{A}^{P}$, $P\geq 1$, where $\mathcal{A}^{P}$ represents the domain of all possible hyperparameter values. For time series data, augmentations could involve mean shifts, trend injections, or other transformations relevant to the time-series domain, with $\mathcal{A}^{P}$ encompassing the respective hyperparameter space for a given transformation. We describe augmentation in more detail in Sec. 3.1.
2.4 Wasserstein Distance
The Wasserstein distance [1], a distance measure between probability distributions, of order $p$ between any two marginal distributions $\mu$ and $\rho$ is given as
[…]
@@ 3 TSAP: Time Series Anomalies on AutoPilot
Figure 1: Our TSAP framework for end-to-end self-tuning TSAD. Left: Offline trained, differentiable augmentation model $f_{\mathrm{aug}}(\cdot;\boldsymbol{\phi})$ takes as input the normal data and augmentation hyperparameter(s) $\mathbf{a}$, and outputs pseudo-anomalies $\mathbf{\tilde{x}}_{\mathrm{aug}}$. Right: Self-tuning engine incorporates the pre-trained $f_{\mathrm{aug}}$ (with parameters $\boldsymbol{\phi}$ frozen), alternating between two phases: ($i$) detection phase – given $\mathbf{a}^{(t)}$ at iteration $t$, estimate parameters $\boldsymbol{\theta}^{(t)}$ of detector $f_{\mathrm{det}}$ (consisting of Encoderθ and discriminator MLPθ), by optimizing $\mathcal{L}_{\mathrm{trn}}$ (classification loss); ($ii$) alignment phase – given $f_{\mathrm{det}}^{\mathrm{enc}}(\cdot;\boldsymbol{\theta}^{(t)})$, update augmentation (governed by $\mathbf{a}$) to better align the embedded time series $\mathbf{z}_{\mathrm{trn}}\cup\mathbf{z}_{\mathrm{aug}}$ with $\mathbf{z}_{\mathrm{val}}$ in the learned discriminative space. Note that $\mathbf{x}_{\mathrm{val}}$ contains both normal and anomalous time series, but labels are not known or used at any point during training time.
@@
  • C2.
    Comparable Validation Loss: Formulating a validation loss that effectively quantifies alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{test}}$ while being comparable across different hyperparameter initializations.
To address C1 and C2, we center our framework around two main ideas:
Main ideas
  • I1.
    Differentiable Augmentation Module: TSAP implements augmentation function as an Encoder-Decoder neural network, $f_{\mathrm{aug}}$ parameterized by $\boldsymbol{\phi}$, capable of approximating the anomaly-generating mechanism conditioned on $\mathbf{a}\in\mathcal{A}^{P}$. Importantly, this module is pre-trained independently and prior to the initiation of I2, establishing it as an offline component of the framework.
  • I2.
    Self-Tuning Module: At test time online, TSAP iteratively refines the detector $f_{\mathrm{det}}$’s parameters $\boldsymbol{\theta}$ as well as augmentation hyperparameters $\mathbf{a}$, through alternating detection and alignment phases. Alignment is performed on part of the unlabeled $\mathcal{D}_{\mathrm{test}}$, referred to as $\mathcal{D}_{\mathrm{val}}$.
Based on the key ideas I1 and I2, we propose TSAP, a self-tuning self-supervised TSAD framework illustrated in Fig. 1.
Figure 2: Examples of six different types of time series anomalies; (black) original real-world time series, (red) pseudo anomalies generated by $g$. The goal for $f_{\mathrm{aug}}$ is to learn to inject anomalies into inliers, conditional on the discrete (anomaly type) as well as continuous hyperparameters such as level (magnitude), location (start point) and length (see bottom right). More details are presented in Sec. 3.1 and Appx. 0.A.1.
@@ 3.1 Differentiable Augmentation Module
We carefully consider accommodating six types of anomalies that are common in real-world time series; namely, trend, extremum, amplitude, mean shift, frequency shift, and platform. We provide a detailed description of the anomalies in Appx. 0.A.1 and for brevity, illustrate them by examples in Fig. 2 with red lines. Each type of anomaly has three hyperparameters; including its starting position (location), duration (length), and severity (level) as illustrated in Fig. 2 (bottom right). Extremum captures a spike and only has two hyperparameters as its duration is always $1$. Based on $\mathcal{D}_{\mathrm{trn}}$, the anomaly generation scheme $g$ creates an augmented dataset $\mathcal{D}_{\mathrm{aug}}=\{g(\mathbf{x}_{\mathrm{trn}};\mathbf{a})\mid\mathbf{x}_{\mathrm{trn}}\in\mathcal{D}_{\mathrm{trn}},\mathbf{a}\sim\mathcal{A}^{P}\}$, where $\mathbf{a}$ is the vector of augmentation hyperparameters that are uniformly randomly sampled from the domain $\mathcal{A}^{P}$.
@@
To build an augmentation model $f_{\mathrm{aug}}$ for time series, first we use a convolutional neural network (CNN) to encode or embed the input time series $\mathbf{x}_{\mathrm{trn}}$ into the feature map $\mathbf{z}_{\mathrm{trn}}$. We then encode the augmentation hyperparameters $\mathbf{a}$ into $\mathbf{z}_{\mathrm{a}}$, which has the same shape as $\mathbf{z}_{\mathrm{trn}}$, with a multilayer perceptron (MLP). Since the feature map $\mathbf{z}_{\mathrm{trn}}$ generated by the CNN encoder keeps the positional information of the original time series, adding $\mathbf{z}_{\mathrm{a}}$ to $\mathbf{z}_{\mathrm{trn}}$ ensures that only the part with the desired location and length in $\mathbf{z}_{\mathrm{aug}}$ is manipulated. To ensure that the feature maps $\mathbf{z}_{\mathrm{aug}}$ and $\mathbf{z}_{\mathrm{trn}}$ are in the same embedding space, they share the same decoder to reconstruct back to the time series $\mathbf{x}_{\mathrm{aug}}=g(\mathbf{x}_{\mathrm{trn}};\mathbf{a})$ and $\mathbf{x}_{\mathrm{trn}}$, respectively. As such, the loss function of $f_{\mathrm{aug}}$ is based on the reconstruction of both $\mathcal{D}_{\mathrm{trn}}$ and $\mathcal{D}_{\mathrm{aug}}$:
@@
where $\mathbf{\tilde{x}}_{\mathrm{aug}}$ denotes the output of $f_{\mathrm{aug}}(\mathbf{x}_{\mathrm{trn}};\mathbf{a})$ for the same value of the hyperparameters sampled by $g$ (see Fig. 1 left).
@@ 3.2 Self-Tuning Module
Phase ($i$): Detection
This phase focuses on estimating the parameters $\boldsymbol{\theta}^{(t)}$ of the detector $f_{\mathrm{det}}$ (comprising of an encoder $f_{\mathrm{det}}^{\mathrm{enc}}$ and a discriminator $f_{\mathrm{det}}^{\mathrm{mlp}}$) by minimizing the cross-entropy loss $\mathcal{L}_{\mathrm{trn}}$. This aims to classify between the normal samples $\mathbf{x}_{\mathrm{trn}}$ and the augmented pseudo anomalies $\mathbf{\tilde{x}}_{\mathrm{aug}}$ by their embeddings $\mathcal{Z}_{\mathrm{trn}}$ and $\mathcal{Z}_{\mathrm{aug}}$, where $\mathcal{Z}_{\mathrm{trn}}=\{f_{\mathrm{det}}^{\mathrm{enc}}(\mathbf{x})\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$ and $\mathcal{Z}_{\mathrm{aug}}=\{f_{\mathrm{det}}^{\mathrm{enc}}(f_{\mathrm{aug}}(\mathbf{x};\mathbf{a}^{(t)}))\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$ denote the embeddings of the training data and augmented data, respectively, given the current fixed $\mathbf{a}^{(t)}$ at iteration $t$. Note that the parameters $\boldsymbol{\phi}$ of the augmentation model $f_{\mathrm{aug}}$ are frozen throughout this phase.
Phase ($ii$): Alignment
Subsequently, the alignment phase adjusts $\mathbf{a}$ to optimize the unsupervised differentiable validation loss $\mathcal{L}_{\mathrm{val}}$, computed based on the embeddings from the now-updated $f_{\mathrm{det}}^{\mathrm{enc}}$. $\mathcal{L}_{\mathrm{val}}$’s objective is to measure the degree of alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$ in the embedding space, as expressed by the Wasserstein distance in Eq. (1). This metric is chosen for its effectiveness in capturing the overall distributional discrepancies between datasets, offering a more complete comparison than mere point-wise metrics [1], which is especially important in TSAD given the often subtle nature of time series anomalies. Note that the embeddings are normalized to ensure scale invariance before being passed to $\mathcal{L}_{\mathrm{val}}$, so as to avoid the trivial solution of achieving alignment by setting all embeddings to the zero vector [34]. Lastly, as the embeddings $\{\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}}\}$ in $\mathcal{L}_{\mathrm{val}}$ are obtained through the updated $f_{\mathrm{det}}^{\mathrm{enc}}$, the optimization process needs to track the change in $\boldsymbol{\theta}$ caused by the update of $\mathbf{a}$. As such, TSAP employs a second-order optimization process, similar to [34].
@@
While Algo. 1 describes continuous hyperparameter tuning, the discrete hyperparameter (i.e. augmentation type) selection, corresponding to a specific anomaly type, is done through grid search as the number of anomaly types is finite. Hence, we initialize TSAP for different augmentation types and compare $\mathcal{L}_{\mathrm{val}}$ across types to select the one that yields the best alignment. The idea is that the wrong augmentation type will have poor alignment, while one that reflects the true anomalies in $\mathcal{D}_{\mathrm{test}}$ will result in better alignment, granted proper tuning of the continuous hyperparameters through Algo. 1.

### v2

@@ Abstract
Time series anomaly detection (TSAD) finds many applications such as monitoring environmental sensors, industry KPIs, patient biomarkers, etc. A two-fold challenge for TSAD is a versatile and unsupervised model that can detect various different types of time series anomalies (spikes, discontinuities, trend shifts, etc.) without any labeled data. Self-supervised models in particular tackle unsupervised TSAD by transforming the input via various augmentations to create pseudo anomalies for training. However, their performance is sensitive to the choice of augmentation, which is hard to choose in practice, while there exists no effort in the literature on data augmentation tuning for TSAD without labels. Our work aims to fill this gap. We introduce TSAP for TSA “on autoPilot”, which can (self-)tune augmentation hyperparameters end-to-end. It stands on two key components: a differentiable augmentation architecture and an unsupervised validation loss to effectively assess the alignment between augmentation type and anomaly type. Case studies show TSAP’s ability to select the augmentation type and associated hyperparameters.
@@ 1 Introduction
Anomaly detection (AD) is a critical task in various domains such as cybersecurity, healthcare, and finance. AD is especially important in time series to ensure system safety and reliability. Thus, there exists a large body of work on time series AD as presented in various surveys [1, 2]. Recent advances in self-supervised learning (SSL) have significantly impacted AD by surpassing traditional unsupervised (or one-class) learning approaches. SSL’s key advantage lies in its ability to self-generate labeled samples, or pseudo anomalies, enabling a more focused exploration of a plausible subspace rather than an exhaustive, impractical search of the entire space. Central to SSL-based AD are data augmentation functions, which create such pseudo labels. These are then used for the (self-)supervised training of an anomaly detector, such as by predicting whether the input is augmented or not [3], which augmentation function is used [4], or through contrastive learning [5]. The success of these approaches highly depends on how well the augmented data mimics the true anomalies [6].
In this paper we introduce TSAP, a novel approach for SSL-based time series anomaly detection (TSAD) “on autoPilot” equipped with end-to-end augmentation hyperparameter tuning. Tuning both discrete and continuous hyperparameters of augmentation enables TSAP to be an anomaly detector that is most suitable for a given task. We summarize our main contributions as follows:
@@
    Problem: Our work is the first attempt to tune both discrete and continuous hyperparameters of data augmentation in SSL-based TSAD, without labels at training time.
@@
    New TSAD Method: We propose TSAP  [^1] Code and datasets available at: https://github.com/B-Deforce/TSA-on-autoPilot., which accommodates various time series anomaly types and enables automatic tuning of related hyperparameters (e.g., magnitude, duration) with a differentiable validation loss quantifying alignment between augmented and unlabeled test data.
@@
    Effectiveness: By carefully selecting augmentation type and its (continuous) hyperparameters, TSAP outperforms existing unsupervised and self-supervised approaches, including the SOTA NeuTraL-AD [7] which also employs learnable augmentations.
@@ 2 TSAP: Time Series Anomalies on AutoPilot
Figure 1: Our TSAP framework for end-to-end self-tuning TSAD. Left: Offline trained, differentiable augmentation model $f_{\mathrm{aug}}(\cdot;\boldsymbol{\phi})$ takes as input the normal data and augmentation hyperparameter(s) $\mathbf{a}$, and outputs pseudo-anomalies $\mathbf{\tilde{x}}_{\mathrm{aug}}$. Right: Self-tuning engine incorporates the pre-trained $f_{\mathrm{aug}}$ (with parameters $\boldsymbol{\phi}$ frozen), alternating between two phases as outlined in Algo. 1.
@@
  2. C2.
    Comparable Validation Loss: Formulating a validation loss that quantifies alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{test}}$ while being comparable across different hyperparameter initializations.
To address C1 and C2, we center our framework around two main ideas. First, TSAP employs a differentiable augmentation module implemented as an Encoder-Decoder neural network, $f_{\mathrm{aug}}$, parameterized by $\boldsymbol{\phi}$. This module approximates the anomaly-generating mechanism, conditioned on $\mathbf{a}\in\mathcal{A}^{P}$, $P\geq 1$, where $\mathcal{A}^{P}$ represents the domain of all possible hyperparameter values (e.g., magnitude, duration, …). Importantly, this module is pre-trained independently, establishing it as an offline component of the framework. Second, at test time online, TSAP iteratively refines the detector $f_{\mathrm{det}}$’s parameters $\boldsymbol{\theta}$ as well as augmentation hyperparameters $\mathbf{a}$, through alternating detection and alignment phases. Alignment is performed on part of the unlabeled $\mathcal{D}_{\mathrm{test}}$, referred to as $\mathcal{D}_{\mathrm{val}}$. Based on this, we propose TSAP, a self-tuning self-supervised TSAD framework illustrated in Fig. 1.
@@ 2.1 Differentiable Augmentation Module
We carefully consider accommodating six types of anomalies that are common in real-world time series; namely, trend, extremum, amplitude, mean shift, frequency shift, and platform. We provide a detailed description with visual examples of the anomalies in Appx. A. Each type of anomaly has three hyperparameters; including its starting position (location), duration (length), and severity (level). Extremum captures a spike and only has two hyperparameters as its duration is always $1$. Based on $\mathcal{D}_{\mathrm{trn}}$, the anomaly generation scheme $g$ creates an augmented dataset $\mathcal{D}_{\mathrm{aug}}=\{g(\mathbf{x}_{\mathrm{trn}};\mathbf{a})\mid\mathbf{x}_{\mathrm{trn}}\in\mathcal{D}_{\mathrm{trn}},\mathbf{a}\sim\mathcal{A}^{P}\}$, where $\mathbf{a}$ is the vector of augmentation hyperparameters that are uniformly randomly sampled from the hyperparameter domain $\mathcal{A}^{P}$.
@@
To build an augmentation model $f_{\mathrm{aug}}$ for time series, we use a convolutional neural network (CNN) to encode the input time series $\mathbf{x}_{\mathrm{trn}}$ into the feature map $\mathbf{z}_{\mathrm{trn}}$. We then encode the augmentation hyperparameters $\mathbf{a}$ into $\mathbf{z}_{\mathrm{a}}$, which has the same shape as $\mathbf{z}_{\mathrm{trn}}$, with a multilayer perceptron (MLP). Since the feature map $\mathbf{z}_{\mathrm{trn}}$ generated by the CNN encoder keeps the positional information of the original time series, adding $\mathbf{z}_{\mathrm{a}}$ to $\mathbf{z}_{\mathrm{trn}}$ ensures that only the part with the desired location and length in $\mathbf{z}_{\mathrm{aug}}$ is manipulated. To ensure that the feature maps $\mathbf{z}_{\mathrm{aug}}$ and $\mathbf{z}_{\mathrm{trn}}$ are in the same embedding space, they share the same decoder to reconstruct back to the time series $\mathbf{x}_{\mathrm{aug}}=g(\mathbf{x}_{\mathrm{trn}};\mathbf{a})$ and $\mathbf{x}_{\mathrm{trn}}$. As such, the loss function of $f_{\mathrm{aug}}$ is based on the reconstruction of both $\mathcal{D}_{\mathrm{trn}}$ and $\mathcal{D}_{\mathrm{aug}}$:
@@
where $\mathbf{\tilde{x}}_{\mathrm{aug}}$ denotes the output of $f_{\mathrm{aug}}(\mathbf{x}_{\mathrm{trn}};\mathbf{a})$ for the same $\mathbf{a}$ as sampled by $g$ (see Fig. 1 left).
2.2 Self-Tuning Module
@@
Phase ($i$) Detection
This phase is focused on estimating the parameters $\boldsymbol{\theta}^{(t)}$ of the detector $f_{\mathrm{det}}$ (comprising of an encoder $f_{\mathrm{det}}^{\mathrm{enc}}$ and a discriminator $f_{\mathrm{det}}^{\mathrm{mlp}}$) by minimizing the cross-entropy loss $\mathcal{L}_{\mathrm{trn}}$. This aims to classify between the normal samples $\mathbf{x}_{\mathrm{trn}}$ and the augmented pseudo anomalies $\mathbf{\tilde{x}}_{\mathrm{aug}}$ by their embeddings $\mathcal{Z}_{\mathrm{trn}}$ and $\mathcal{Z}_{\mathrm{aug}}$, where $\mathcal{Z}_{\mathrm{trn}}=\{f_{\mathrm{det}}^{\mathrm{enc}}(\mathbf{x})\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$ and $\mathcal{Z}_{\mathrm{aug}}=\{f_{\mathrm{det}}^{\mathrm{enc}}(f_{\mathrm{aug}}(\mathbf{x};\mathbf{a}^{(t)}))\mid\mathbf{x}\in\mathcal{D}_{\mathrm{trn}}\}$ denote the embeddings of the training data and augmented data, given the current fixed $\mathbf{a}^{(t)}$ at iteration $t$. Note that the parameters $\boldsymbol{\phi}$ of $f_{\mathrm{aug}}$ are frozen throughout this phase.
Phase ($ii$) Alignment
Subsequently, the alignment phase adjusts $\mathbf{a}$ to optimize the unsupervised differentiable validation loss $\mathcal{L}_{\mathrm{val}}$, computed based on the embeddings from the now-updated $f_{\mathrm{det}}^{\mathrm{enc}}$. $\mathcal{L}_{\mathrm{val}}$’s objective is to measure the degree of alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$ in the embedding space, as expressed by the Wasserstein distance in Eq. (2), Appx. B.1. Note that the embeddings are normalized to ensure scale invariance before being passed to $\mathcal{L}_{\mathrm{val}}$. This avoids the trivial solution of achieving alignment by setting all embeddings to $\mathbf{0}$ [8]. As the embeddings $\{\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}}\}$ in $\mathcal{L}_{\mathrm{val}}$ are obtained through the updated $f_{\mathrm{det}}^{\mathrm{enc}}$, the optimization needs to track the change in $\boldsymbol{\theta}$ caused by the update of $\mathbf{a}$. As such, TSAP employs a second-order optimization process, similar to [8].
@@
While Algo. 1 describes continuous hyperparameter tuning, the discrete hyperparameter (i.e. augmentation/anomaly type) selection is done through grid search as the number of anomaly types is finite. Hence, we initialize TSAP for different augmentation types and compare $\mathcal{L}_{\mathrm{val}}$ across types to select the one that yields the best alignment. The idea is that the wrong augmentation type will have poor alignment, while one that reflects the true anomalies in $\mathcal{D}_{\mathrm{test}}$ will result in better alignment, granted proper tuning of the continuous hyperparameters through Algo. 1.